OpenStudy (camerondoherty):

Can someone show me how to complete the square of: \[f(t) = -16t^2 - 32t + 128\]

OpenStudy (mathmath333):

first thing u should do is take out common any coefficients if possible

OpenStudy (camerondoherty):

I think you have to set f(t) to 0 right?

OpenStudy (just_one_last_goodbye):

hey do you got this one Al Capone?

OpenStudy (camerondoherty):

So um, \[0=16t^2+32t+c?\]

OpenStudy (camerondoherty):

Then don't you have to find a c value that completes the square?

OpenStudy (just_one_last_goodbye):

Hey Cammi still need help? ^-^

OpenStudy (camerondoherty):

I do

OpenStudy (just_one_last_goodbye):

sorry froze ok so..... lets see...

OpenStudy (just_one_last_goodbye):

\[0 = 16t _{2} + 32t + c\]

OpenStudy (camerondoherty):

You mean t^2 right?

OpenStudy (just_one_last_goodbye):

I think thats right ._. and no Idk but im getting the answer with 16t_(2)

OpenStudy (camerondoherty):

Thats not the soution though right? cx Im completing the square, not factoring

OpenStudy (just_one_last_goodbye):

oh xD sorry wrong steps

OpenStudy (camerondoherty):

I'll just do it laterz Thanx for trying though cx

OpenStudy (just_one_last_goodbye):

hey ill start solving it if you want c: so when "laterz" comes i can have it ready

OpenStudy (mathmath333):

why did u replace \(c\) with \(128\)

OpenStudy (triciaal):

f(t)=−16t2−32t+128

OpenStudy (camerondoherty):

Thats how I remember it from my textbook I might've interperted it wrongly

OpenStudy (triciaal):

@camerondoherty you do not change what is given you set the expression equal to 0

OpenStudy (triciaal):

so -16t^2 - 32 t + 128 = 0

OpenStudy (camerondoherty):

So \[0=16t^2+32t+128 \]

OpenStudy (triciaal):

as suggested by someone already since you see a common factor you can factor that out

OpenStudy (triciaal):

@mathmath333 will continue

OpenStudy (camerondoherty):

Yes, but the question asks to complete the square. I've already factored it cx

OpenStudy (mathmath333):

see if this makes sense \(\large\tt \begin{align} \color{black}{f(t) = -16(t^2 + 2t -8)\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (mathmath333):

this \(\Huge \uparrow \) makes it easy

OpenStudy (camerondoherty):

But I have already factored it out, I just wanted to find the vertex by completing the square

OpenStudy (mathmath333):

yes m telling the method of completing the square we know that \(\large\tt \begin{align} \color{black}{t^2 + 2t+1=(t+1)^2\hspace{.33em}\\~\\}\end{align}\) so we just adjust that in \(\large\tt \begin{align} \color{black}{f(t) = -16(t^2 + 2t -8)\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (mathmath333):

this can be done by writting it as \(\large\tt \begin{align} \color{black}{f(t) = -16(t^2 + 2t \color{red}{+1-1}-8)\hspace{.33em}\\~\\}\end{align}\) see if it makes sense

OpenStudy (camerondoherty):

Not really >.< Sorry...

OpenStudy (camerondoherty):

I have to go too... sorry cx

OpenStudy (mathmath333):

we just want to put this \(\large\tt \begin{align} \color{black}{t^2 + 2t+1=(t+1)^2\hspace{.33em}\\~\\}\end{align}\)